OpenStudy (anonymous):

y=5/8x + 5 a) -5/8 b) 5/8 c) -5 d) 5

OpenStudy (johnweldon1993):

what exactly is the question asking??

OpenStudy (anonymous):

identify the y-intercept.

OpenStudy (johnweldon1993):

Ahh okay....well in the form y = mx + b m is the slope and b is the y intercept so y = 5/8x + 5 y = m x + b what is your y-intercept?

OpenStudy (anonymous):

hmmmm 5?

OpenStudy (johnweldon1993):

correct!

OpenStudy (anonymous):

sweet thx for the help and i am prob going to need more help haha